In an ap, the sum of the first n terms is 5n^2/2+3n/2. Find its 20th terms

Grade:10

1 Answers

Arun
6 years ago

Sn = (5n+ 3n) / 2

Therefore, S20 = [5 (20)+ 3 (20) ] / 2

Thus S20 = 2060 / 2 = 1030

 

S19 = [5 (19)2 + 3 (19)] / 2  = 1862 / 2

Thus, S19 = 931

 

T=  Sn - Sn-1

Therefore, T20 = S20 - S19

T20 =  1030 - 931 = 99